OpenStudy (anonymous):

This question wants me to know which functions have a positive rate of change and which have a negative rate of change? So the answer must be either positive rate or negative rate. Could some one help with four practice questions by chance? XD

OpenStudy (anonymous):

The first one is 2x – 3y = 9 ... I think its a positive rate..? Im sort of guessing XD

OpenStudy (anonymous):

@swissgirl @jim_thompson5910 Could either of you help me??

jimthompson5910 (jim_thompson5910):

are you able to solve 2x – 3y = 9 for y?

OpenStudy (anonymous):

um... i think its... i dont know XD i think i need help with that too

jimthompson5910 (jim_thompson5910):

start by subtracting 2x from both sides

OpenStudy (anonymous):

ok so thats -3y = 9 - 2x

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

that's the same as -3y = -2x + 9

jimthompson5910 (jim_thompson5910):

how do we go from -3y to just y ?

OpenStudy (anonymous):

we divide 3 from both sides?

jimthompson5910 (jim_thompson5910):

-3 actually

jimthompson5910 (jim_thompson5910):

but you have the right idea

OpenStudy (anonymous):

oh yeah oop XD

OpenStudy (anonymous):

so thats y = -2x + -3 ?

jimthompson5910 (jim_thompson5910):

you forgot to divide the -2x by -3

jimthompson5910 (jim_thompson5910):